For a thermodynamic system, isobaric coefficient of volume expansion (alpa) and isothermal compressibility (bita) are defined as

ailpa=1/V(dV/dT)p

bita=-1/V(dV/dT)T

Show that for an isochoric change, dp = dT.



1
Expert's answer
2021-03-09T15:29:47-0500

Given,

"\alpha = \frac{1}{V}(\frac{dV}{dT})_p"


"\beta = - \frac{1}{V}(\frac{dV}{dP})_T"


Now, from the ideal gas equation,


"PV=nRT"


"V=\frac{nRT}{P}"


Now, taking the differentiation with respect to the corresponding terms,


"\frac{dV}{dT}=\frac{nR}{P}"


if P is constant then,


"(\frac{dV}{dT})_P=\frac{nR}{P}...(i)"


now, taking the differentiation with respect to P,

"(\frac{dV}{dP})_T=\frac{nRT}{P^2}...(ii)"

Now, taking the ratio of "(i)" and "(ii)"


"\frac{(\frac{dV}{dT})_P}{(\frac{dV}{dP})_T}=\frac{\frac{nR}{P}}{\frac{nRT}{P^2}}"


"\frac{(dP)_T}{(dT)_P}=\frac{P}{T}"


"(dP)_T=(dT)_P\times \frac{P}{T}"